a particle moves on a curve of radius=2m with a speed v=t2 find total acceleration at 4 second
 
 
plz explain everything indetail
 

Grade:11

1 Answers

Vikas TU
7 years ago
Tagential accln. =>
at = dv/dt => 2t at t= 4s  
at = 2*4 = 8 m/s^2
ac = v^2/r = 256/2 = 128 m/s^2
Resulatant = root(ac^2 + at^2)